I don't understand it says to solve by the addition or subtraction method
2x+y=18
2x-y=22

Answer by Mathtut(3670)

if you add these 2 equations together you would be adding the x terms to x terms the y terms to the y terms and the constants to the constants.
so 2x+2x=4x--->y+(-y)=0 and 18+22=40
so your left with 4x=40 or x=10
now plug x's value back into either of the original equations to find y
I choose the 1st 2(10)+y=18 subtract 20 from either side and you get y=-2